# Multiple Imputation for Nonresponse in Complex Surveys Using Design Weights and Auxiliary Margins

### v4

@@ 1 Introduction
In this short note, we consider settings where analysts seek to integrate auxiliary information on marginal distributions in multiple imputation for nonresponse (Rubin, 1987) in complex surveys. We build on the missing data with auxiliary margins, or MD-AM, framework introduced by Akande et al. (2021) and extended to complex surveys by Akande and Reiter (2022), Tang et al. (2024), and Yang and Reiter (2025). The latter three works impute missing values to ensure completed-data, survey-weighted inferences are plausible given the known margins. However, these works impose a simplifying condition on the unit nonrespondents, namely that the analyst does not use their design weights and instead replaces these weights with a convenient constant. This creates difficulties for design-based analyses, as the resulting weights for the full sample do not correspond to the design.
We propose MD-AM models that allow use of the design-based weights for all sampled units. The basic idea is as follows. We first impute all missing values due to item nonresponse using multiple imputation by chained equations (Van Buuren, 2018; Raghunathan et al., 2001). In each partially completed dataset, we then impute the unit nonrespondents’ values of the categorical variables that have auxiliary margins. To do so, we specify probabilities for imputations so that, in expectation, imputations result in plausible Horvitz and Thompson (1952) estimates in the completed data given the known margins, using the design weights for all sampled units. For the variables without auxiliary margins, we use the random hot deck imputation proposed in Yang and Reiter (2025).
@@
The remainder of this note is organized as follows. Section 2 provides a brief review of the hybrid missing MD-AM model (Tang et al., 2024; Yang and Reiter, 2025), which we extend to incorporate design weights for unit nonrespondents. Section 3 introduces the methodology. Section 4 illustrates the methods using simulation studies. Codes are available at https://github.com/kevinxu47/MDAM.
2 Hybrid Missingness MD-AM Modeling
@@
In reviewing the hybrid missingness MD-AM model, we closely follow the notation in Yang and Reiter (2025). Let $\mathcal{P}$ be a finite population comprising $N$ units. For $i=1,\dots,N$, let $\boldsymbol{x}_{i}=(x_{i1},\dots,x_{ip})$ denote the $p$ survey variables for unit $i$; let $I_{i}=1$ if unit $i$ is selected for inclusion in the survey and $I_{i}=0$ otherwise; let $\pi_{i}=\Pr(I_{i}=1)$ represent the probability that unit $i$ is selected into the survey; and, let $w_{i}^{d}=1/\pi_{i}$ be its design weight. Let $n=\sum_{i=1}^{N}I_{i}$ be the intended sample size of the survey. We refer to the sampled units as $\mathcal{S}$.
@@
Let $U$ be the unit nonresponse indicator so that, for each unit $i\in\mathcal{S}$, $U_{i}=1$ if the unit provides no responses to any survey questions and $U_{i}=0$ otherwise. Let $\boldsymbol{R}=(R_{1},\dots,R_{p})$ be item nonresponse indicators corresponding to $\boldsymbol{X}=(X_{1},\dots,X_{p})$ so that, for any unit $i$ with $U_{i}=0$ and any variable $X_{j}$, $R_{ij}=1$ if unit $i$ does not respond to question $X_{j}$ and $R_{ij}=0$ otherwise. When $U_{i}=1$, $(R_{i1},\dots,R_{ij})$ are undefined.
@@
The survey organization has auxiliary information $\mathcal{A}$ on the marginal distributions of a subset of the $p$ survey variables. For $j=1,\dots,p$, let $\mathcal{A}_{j}$ comprise the known marginal for $X_{j}$ when it exists. We write $X_{j}\in\mathcal{A}$ whenever $X_{j}$ has a known margin in $\mathcal{A}$. For convenience, we presume each $\mathcal{A}_{j}$ comprises population totals; incorporating percentages is a simple modification. For any categorical $X_{j}$ taking on levels $c=\{1,\dots,m_{j}\}$ and $X_{j}\in\mathcal{A}$, let $T_{jc}=\sum_{i=1}^{N}I(x_{ij}=c)$ be the total number of units in $\mathcal{P}$ at level $c$. Here, $I(\cdot)=1$ when the condition inside the parenthesis is true and $I(\cdot)=0$ otherwise.
@@
The hybrid missingness MD-AM model specifies a joint distribution of $(\boldsymbol{X},\boldsymbol{R},U)$. It uses a pattern mixture model for unit nonresponse coupled with selection models for item nonresponse. Specifically, the model presumes $U\sim\textrm{Bernoulli}(\pi_{u})$, where $\pi_{u}=\Pr(U=1)$ is the marginal probability of unit nonresponse. For any $j=1,\dots,p$, let $g_{j}(-)$ represent the model for $X_{j}$ given $X_{1},\dots,X_{j-1}$, and let $\Omega_{j}$ and $\theta_{j}$ be the corresponding model parameters. For example, $g_{j}(-)$ could be a logistic regression of $X_{j}$ on some function of $X_{1},\dots,X_{j-1}$, possibly including interactions, as well as a main effect for $U$ when $X_{j}\in\mathcal{A}$. Thus, for $\boldsymbol{X}|U$, the hybrid missingness MD-AM model uses
@@
The imputation and estimation algorithm in Tang et al. (2024) does not use the design weights for the unit nonrespondents. Instead, they replace the design weights for unit nonrespondents with a constant that makes $\sum_{i=1}^{N}w_{i}=N$; that is, they base imputations and completed-data inferences on, for all units $i\in\mathcal{S},$
@@
For any unit $i\in\mathcal{S}$ and $j=1,\dots,p$, let $x_{ij}^{\star}=x_{ij}$ when $(R_{ij}=0,U_{i}=0)$, and let $x_{ij}^{\star}$ be an imputed value when $R_{ij}=1$ or $U_{i}=1$. For any $X_{j}\in\mathcal{A}$, Akande and Reiter (2022), Tang et al. (2024), and Yang and Reiter (2025) suppose the imputations to adhere to (2.1) – (2.3), with the additional constraint that
$\hat{T}_{jc}^{*}=\sum_{i\in\mathcal{S}}w_{i}I(x_{ij}^{\star}=c)\sim N(T_{jc},V_{jc}).$ (2.5)
Here, $V_{jc}$ is set by the analyst and determines how closely $\hat{T}_{jc}^{*}$ matches $T_{jc}$ in any completed dataset. As we also assume (2.5), we discuss setting $V_{jc}$ in Section 3.
@@ 3 Methodology
We now modify the hybrid missingness MD-AM model to use the design weights for all sampled individuals, including unit nonrespondents. For convenience, we presume that $X_{j}\in\mathcal{A}$ for $j=1,\dots,k<p$, and the remaining $p-k$ variables do not have auxiliary margins.
@@ 3.1 Variables with Item Nonresponse
We first create multiple imputations for all values of $x_{ij}$ missing due to item nonresponse for units with $U_{i}=0$. Following Yang and Reiter (2025), we implement multiple imputation by chained equations (MICE) using only units with $U_{i}=0$. The specification of the MICE algorithm does not require any extraordinary considerations. In our simulations, we use the “mice” package in R (Van Buuren, 2018), following its default settings including the ordering of variables. Each conditional model is a function of all other survey variables except $X_{j}$, including possibly interaction terms. The conditional models also could include subsets of $(R_{1},\dots,R_{j-1},R_{j+1},\dots,R_{p})$ as predictors; these are fully observed for units with $U_{i}=0$. We run the MICE procedure to create $L$ completed datasets for the units with $U_{i}=0$.
@@ 3.2 Variables with Margins for Unit Nonrespondents
After imputing values for item nonresponse, in each completed dataset we impute the unit nonrespondents’ values for all $X_{j}\in\mathcal{A}$. The analyst orders these variables based on ease of modeling. For notational convenience, we suppose that $X_{1}$ is imputed first, $X_{2}$ is imputed second, and so on sequentially until $X_{k}.$
To begin, for each unit $i$ with $U_{i}=1$ and in each of the $L$ completed datasets, the analyst sets an initial probability distribution for imputing $x_{i1}$. For example, in each completed dataset, the analyst estimates a (multinomial) logistic regression of $X_{1}$ on some function of $W$ and uses the predicted probabilities derived from the estimated model. Alternatively, the analyst could regress $X_{1}$ on an intercept only or compute marginal probabilities of $X_{1}$ using survey-weighted ratio estimators. For these two alternatives, the initial probabilities are identical for all unit nonrespondents. This approach may be especially appropriate when the analyst does not have strong reason to expect associations between the design weights and the missing values of $X_{1}$ for unit nonrespondents. Regardless of how it is derived, we refer to the initial distribution for imputing $x_{i1}$ as its working distribution, notated as $\{p_{i1c}:c=1,\dots,m_{1};U_{i}=1;i\in\mathcal{S}\}.$ Because of imputation for item nonresponse, the working distribution for any unit can differ across completed datasets; however, for convenience, we forego notation designating completed datasets. The analyst performs the computations of this section for each of the $L$ completed datasets.
Imputing each missing $x_{i1}$ using its working distribution does not utilize the information in $\mathcal{A}_{1}$. Thus, the resulting imputations could generate completed-data Horvitz and Thompson (1952) estimates that are far from the known totals for some levels of $X_{1}$. We therefore modify $\{p_{i1c}\}$ to ensure the resulting imputations result in reasonable completed-data point estimates of these totals.
To do so, we first simulate plausible values of $\hat{T}_{1c}$ for each level $c$ by appealing to large-sample central limit theorems. For $c=1,\dots,(m_{1}-1)$, we sample $\hat{T}_{1c}$ from $N(T_{1c},V_{1c})$ and set $\hat{T}_{1m_{1}}=N-\sum_{c=1}^{m_{1}-1}\hat{T}_{1c}$. We then find modified probabilities $\{\tilde{p}_{i1c}:c=1,\dots,m_{1};U_{i}=1;i\in\mathcal{S}\}$ so that, for $c=1,\dots,m_{1}$, the expectation of each completed-data estimator $\hat{T}_{1c}^{\star}$ from (2.5) over imputations for unit nonresponse approximately equals $\hat{T}_{1c}$. Put another way, recognizing that
$E\left(\sum_{i\in\mathcal{S}}w_{i}I(x_{ij}^{\star}=c)\right)$ $=$ $\sum_{i\in\mathcal{S}}w_{i}I(x_{ij}^{\star}=c)I(U_{i}=0)+E(\sum_{i\in\mathcal{S}}w_{i}I(x_{ij}^{\star}=c)I(U_{i}=1))$
@@
We impute $x_{i1}^{*}$ for each unit with $U_{i}=1$ using a random draw from the adjusted probability mass function, $\{\tilde{p}_{i1c}:c=1,\dots,m_{1}\}$, where $\tilde{p}_{i1m_{1}}=1-\sum_{c=1}^{m_{1}-1}f_{1c}{p}_{i1c}$. If $\sum_{c=1}^{m_{1}-1}f_{1c}{p}_{i1c}>1$, we compute (3.1) for $c=1,\dots,m_{1}$ and set $\tilde{p}_{i1c}=f_{1c}{p}_{i1c}/\sum_{c=1}^{m_{1}}f_{1c}{p}_{i1c}$.
When each $p_{i1c}$ equals some common value $p_{1c}$ for all units with $U_{i}=1$, each $\tilde{p}_{i1c}=f_{1c}p_{1c}=\tilde{p}_{1c}$ is given conveniently by
$\tilde{p}_{1c}=\frac{\hat{T}_{1c}-\sum_{i\in\mathcal{S}}w_{i}I(x_{i1}^{*}=c,U_{i}=0)}{\sum_{i\in\mathcal{S}}w_{i}I(U_{i}=1)}.$ (3.2)
We now turn to imputing $X_{2}\in\mathcal{A}$ for unit nonrespondents. We follow a similar strategy: in each completed dataset, start by setting each unit nonrespondent’s working distribution for imputation of $x_{i2}^{*}$ given $x_{i1}^{*}$, then adjust the probabilities to make $(\hat{T}_{21}^{*},\dots,\hat{T}_{2m_{2}}^{*})$ approximately match a sampled value of $(\hat{T}_{21},\dots,\hat{T}_{2m_{2}}).$
Let $p_{i2c}=\Pr(X_{i2}=c|x_{i1}^{*}=d,w_{i})$ be the working probability that $X_{2}=c$ for unit $i$, given its imputed $x_{i1}^{*}=d$ and $w_{i}$. We determine these from a (multinomial) logistic regression of $X_{2}$ on $X_{1}$ and possibly some function of $W$, with coefficients estimated from the completed data for units with $U_{i}=0$. If we disregard $W$, $p_{i2c}=\Pr(X_{2}=c|X_{1}=d)$ for $c=1,\dots,m_{2}$ and $d=1,\dots,m_{1}$; that is, the conditional probability is the same for all units with a common value of $x_{i1}^{*}=d$.
For $c=1,\dots,m_{2}-1$, we sample a plausible value of $\hat{T}_{2c}$ by drawing from a $N(T_{2c},V_{2c})$, setting the estimated total for the final level as $\hat{T}_{2m_{2}}=N-\sum_{c=1}^{m_{2}-1}\hat{T}_{2c}$. Akin to (3.1), we then find suitable constants $f_{2c}>0$ to adjust each $p_{i2c}$. We have

### v5

# Multiple Imputation for Nonresponse in Surveys Using Design Weights and Auxiliary Margins

@@ 1 Introduction
We consider settings where analysts seek to integrate auxiliary information on marginal distributions in multiple imputation for nonresponse (Rubin, 1987) in surveys. We build on the missing data with auxiliary margins, or MD-AM, framework introduced by Akande et al. (2021) and extended to surveys by Akande and Reiter (2022), Tang et al. (2024), and Yang and Reiter (2025). The latter three works impute missing values to ensure completed-data, survey-weighted inferences are plausible given the known margins. However, these works impose a simplifying condition on the unit nonrespondents, namely that the analyst does not use their design weights and instead replaces these weights with a convenient constant. This creates difficulties for design-based analyses, as the resulting weights for the full sample do not correspond to the design.
@@
In this short note, we propose MD-AM models that allow use of the design-based weights for all sampled units. Section 2 provides a brief review of the hybrid missing MD-AM model (Tang et al., 2024; Yang and Reiter, 2025), which we extend to incorporate design weights for unit nonrespondents. Section 3 introduces the methodology. Section 4 illustrates the methods using simulation studies. Codes and additional simulation results are available at https://github.com/kevinxu47/MDAM.
2 Review of Hybrid Missingness MD-AM Modeling
@@
In reviewing the hybrid missingness MD-AM model, we closely follow the notation in Yang and Reiter (2025). Let $\mathcal{P}$ be a finite population comprising $N$ units. For $i=1,\dots,N$, let $\boldsymbol{x}_{i}=(x_{i1},\dots,x_{ip})$ denote the $p$ survey variables for unit $i$; let $I_{i}=1$ if unit $i$ is selected for inclusion in the survey and $I_{i}=0$ otherwise; let $\pi_{i}=\Pr(I_{i}=1)$ represent the probability that unit $i$ is selected into the survey; let $w_{i}^{d}=1/\pi_{i}$ be its design weight; and, let $\boldsymbol{z}_{i}$ be its design variables, e.g., size measures or stratum indicators. Let $n=\sum_{i=1}^{N}I_{i}$ be the intended sample size of the survey. We refer to the sampled units as $\mathcal{S}$.
@@
Let $U$ be the unit nonresponse indicator so that, for each unit $i\in\mathcal{S}$, $U_{i}=1$ if the unit provides no responses to any survey questions and $U_{i}=0$ otherwise. Let $\boldsymbol{R}=(R_{1},\dots,R_{p})$ be item nonresponse indicators corresponding to $\boldsymbol{X}=(X_{1},\dots,X_{p})$ so that, for any unit $i$ with $U_{i}=0$ and any $X_{j}$, $R_{ij}=1$ if unit $i$ does not respond to the question for $X_{j}$ and $R_{ij}=0$ otherwise. When $U_{i}=1$, $(R_{i1},\dots,R_{ij})$ are undefined.
@@
The survey organization has auxiliary information $\mathcal{A}$ on the marginal distributions of a subset of the $p$ survey variables. We write $X_{j}\in\mathcal{A}$ whenever $X_{j}$ has a known margin in $\mathcal{A}$, which we write as $\mathcal{A}_{j}$. For convenience, we presume each $\mathcal{A}_{j}$ comprises population totals; incorporating percentages is a simple modification. For any categorical $X_{j}$ taking on levels $c=\{1,\dots,m_{j}\}$, let $T_{jc}=\sum_{i=1}^{N}I(x_{ij}=c)$ be the total number of units in $\mathcal{P}$ at level $c$. Here, $I(\cdot)=1$ when the condition inside the parenthesis is true and $I(\cdot)=0$ otherwise.
@@
As described in Tang et al. (2024) and Yang and Reiter (2025), the hybrid missingness MD-AM model specifies a joint distribution of $(\boldsymbol{X},\boldsymbol{R},U)$. Specifically, the model presumes $U\sim\textrm{Bernoulli}(\pi_{u})$, where $\pi_{u}=\Pr(U=1)$ is the marginal probability of unit nonresponse. For $j=1,\dots,p$, let $g_{j}(-)$ represent the model for $X_{j}$ given $X_{1},\dots,X_{j-1}$, and let $\Omega_{j}$ and $\theta_{j}$ be the corresponding model parameters. For example, $g_{j}(-)$ could be a logistic regression of $X_{j}$ on some function of $X_{1},\dots,X_{j-1}$, possibly including interactions, as well as a main effect for $U$ when $X_{j}\in\mathcal{A}$. Thus, for $\boldsymbol{X}|U$, the hybrid missingness MD-AM model uses
@@
The imputation and estimation algorithms in Tang et al. (2024) and Yang and Reiter (2025) do not use the design weights for the unit nonrespondents. Instead, they replace the design weights for unit nonrespondents with a constant that makes $\sum_{i=1}^{N}w_{i}=N$; that is, they base imputations and completed-data inferences on, for all units $i\in\mathcal{S},$
@@
For any unit $i\in\mathcal{S}$ and $j=1,\dots,p$, let $x_{ij}^{\star}=x_{ij}$ when $(R_{ij}=0,U_{i}=0)$, and let $x_{ij}^{\star}$ be an imputed value when $R_{ij}=1$ or $U_{i}=1$. For any $X_{j}\in\mathcal{A}$, Akande and Reiter (2022), Tang et al. (2024), and Yang and Reiter (2025) suppose the imputations adhere to (2.1) – (2.3), with the additional constraint that
$\hat{T}_{jc}^{*}=\sum_{i\in\mathcal{S}}w_{i}I(x_{ij}^{\star}=c)\sim N(T_{jc},V_{jc}),$ (2.5)
where the weights are from (2.4). Here, $V_{jc}$ is set by the analyst and determines how closely $\hat{T}_{jc}^{*}$ matches $T_{jc}$ in any completed dataset.
Using (2.5) for imputations has commonalities with balanced random imputation (Chauvet et al., 2011; Chauvet and Do Paco, 2018), although the goals differ. In balanced random imputation, the imputed values are required always to satisfy some enforced constraint, e.g., the survey-weighted mean of the imputed values equals the survey-weighted mean of the observed values. Work on balanced random imputation tends to consider single imputations for item nonresponse, presume values are missing (completely) at random, and set constraints based on observed data quantities. In contrast, the hybrid missingness MD-AM model considers multiple imputation for unit nonresponse that is missing not at random, and it uses auxiliary marginal information to establish distributional constraints on imputations.
Although not done in Tang et al. (2024) or Yang and Reiter (2025), the model for each $X_{j}$ or $R_{j}$ also can include the design variables $\boldsymbol{Z}$ as predictors. Including $\boldsymbol{Z}$ can improve explanatory power, and hence ultimately imputation quality and estimation accuracy, when they are strongly associated with the survey variables or nonresponse indicators (Reiter et al., 2006). Alternatively, the models can include some function of $W$ as a predictor. This is common advice for imputation modeling (e.g., Kim et al., 2006; Quartagno et al., 2020) that is particularly salient when the analyst doing the imputations has the survey weights but not the design variables, e.g., because they are confidential. In the MD-AM models of Tang et al. (2024) and Yang and Reiter (2025), however, every unit nonrespondent has the same value of $w_{i}$ given in (2.4). Thus, including $W$ in the models would not influence the imputation probabilities, and hence not contribute to incorporating the design in the imputation process, for the unit nonrespondents.
@@ 3 Methodology
We now modify the hybrid missingness MD-AM model to use the design weights for all sampled individuals, including unit nonrespondents. For brevity of notation, we now let $W$ refer to the design weights $w_{i}^{d}$ rather than the weights in (2.4). We follow the general strategy outlined in Yang and Reiter (2025). First, we impute values for item nonresponse. Second, we use $\mathcal{A}$ to impute values of variables with margins for the unit nonrespondents. Third, we impute values of the remaining variables for unit nonrespondents. Our innovation is to modify the second step so that imputations and the completed datasets use the design weights rather than (2.4) for the unit nonrespondents. This change requires new imputation algorithms, which we present in Section 3.2. For the first and third steps, we use techniques presented in Yang and Reiter (2025), which we briefly summarize in Section 3.1 and Section 3.3. For convenience, we presume that $X_{j}\in\mathcal{A}$ for $j=1,\dots,k<p$, and the remaining pkp-k variables do not have auxiliary margins.
@@ 3.1 Variables with Item Nonresponse
We first create multiple imputations for all values of $x_{ij}$ missing due to item nonresponse for units with $U_{i}=0$. We implement multiple imputation by chained equations (MICE) using only units with $U_{i}=0$ and potentially including $\boldsymbol{Z}$ as predictors. The specification of the MICE algorithm does not require any extraordinary considerations. In our simulations, we use the “mice” package in R (Van Buuren, 2018), following its default settings including the ordering of variables. We run the MICE procedure to create $L$ completed datasets for the units with $U_{i}=0$.
@@ 3.2 Variables with Margins for Unit Nonrespondents
After imputing values for item nonresponse, in each completed dataset we impute the unit nonrespondents’ values for all $X_{j}\in\mathcal{A}$. The analyst orders these variables based on ease of modeling. For convenience, we suppose that $X_{1}$ is imputed first, $X_{2}$ is imputed second, and so on sequentially until $X_{k}.$
To begin, for each unit $i$ with $U_{i}=1$ and in each of the $L$ completed datasets, the analyst sets an initial probability distribution for imputing $x_{i1}$. For example, in each completed dataset, the analyst estimates a (multinomial) logistic regression of $X_{1}$ on some function of $\boldsymbol{Z}$ (or $W$), and uses the predicted probabilities from the estimated model. Alternatively, the analyst could regress $X_{1}$ on an intercept only or compute marginal probabilities of $X_{1}$ using survey-weighted ratio estimators. For these two alternatives, the initial probabilities are identical for all unit nonrespondents. This approach may be especially appropriate when the analyst does not expect associations between the design weights and the missing values of $X_{1}$ for unit nonrespondents.
Regardless of how it is derived, we refer to the initial distribution for imputing $x_{i1}$ as its working distribution, notated as $\{p_{i1c}:c=1,\dots,m_{1};U_{i}=1;i\in\mathcal{S}\}.$ Because of imputation for item nonresponse, the working distribution for any unit can differ across completed datasets; however, for convenience, we forego notation designating completed datasets. The analyst performs the computations of this section for each of the $L$ completed datasets.
Imputing each missing $x_{i1}$ using its working distribution does not utilize the information in $\mathcal{A}_{1}$. The resulting imputations could generate completed-data Horvitz and Thompson (1952) estimates that are far from the known totals for some levels of $X_{1}$, particularly when values are missing not at random. We therefore modify $\{p_{i1c}\}$ to ensure the resulting imputations result in reasonable design-based, completed-data estimates of these totals. This is the primary objective and motivation of our extension of the MD-AM framework.
To do so, we first simulate plausible values of $\hat{T}_{1c}$ for each level $c$ by appealing to large-sample central limit theorems. For $c=1,\dots,(m_{1}-1)$, we sample $\hat{T}_{1c}$ from $N(T_{1c},V_{1c})$ and set $\hat{T}_{1m_{1}}=N-\sum_{c=1}^{m_{1}-1}\hat{T}_{1c}$. We then find adjusted probabilities $\{\tilde{p}_{i1c}:c=1,\dots,m_{1};U_{i}=1;i\in\mathcal{S}\}$ so that, for $c=1,\dots,m_{1}$, the expectation of each completed-data estimator $\hat{T}_{1c}^{\star}$ from (2.5) over imputations for unit nonresponse approximately equals $\hat{T}_{1c}$. Put another way, recognizing that
$\mathbb{E}\left[\sum_{i\in\mathcal{S}}w_{i}I(x_{ij}^{\star}=c)\right]$ $=$ $\sum_{i\in\mathcal{S}}w_{i}I(x_{ij}^{\star}=c)I(U_{i}=0)+\mathbb{E}\left[\sum_{i\in\mathcal{S}}w_{i}I(x_{ij}^{\star}=c)I(U_{i}=1)\right]$
@@
We impute $x_{i1}^{*}$ for each unit with $U_{i}=1$ by drawing randomly from the adjusted probability distribution, $\{\tilde{p}_{i1c}:c=1,\dots,m_{1}\}$, where $\tilde{p}_{i1m_{1}}=1-\sum_{c=1}^{m_{1}-1}f_{1c}{p}_{i1c}$. If $\sum_{c=1}^{m_{1}-1}f_{1c}{p}_{i1c}>1$, we compute (3.1) for $c=1,\dots,m_{1}$ and set $\tilde{p}_{i1c}=f_{1c}{p}_{i1c}/\sum_{c=1}^{m_{1}}f_{1c}{p}_{i1c}$. In the event that some $f_{1c}p_{i1c}<0$, we set those $\tilde{p}_{i1c}=0$. Analysts may consider reducing $V_{1c}$ in this case.
We now turn to imputing $X_{2}\in\mathcal{A}$ for unit nonrespondents. We follow a similar strategy: in each completed dataset, start by setting each unit nonrespondent’s working distribution for imputation of $x_{i2}^{*}$ given $x_{i1}^{*}$ and $(\boldsymbol{z}_{i},w_{i}$), then adjust the probabilities to make $(\hat{T}_{21}^{*},\dots,\hat{T}_{2m_{2}}^{*})$ approximately match a sampled value of $(\hat{T}_{21},\dots,\hat{T}_{2m_{2}}).$ We present two options for imputing $X_{2}$, one based on applying a multiplicative adjustment to the working probabilities (Section 3.2.1) and the other based on solving a system of equations making use of $\mathcal{A}$ (Section 3.2.2).
3.2.1 Multiplicative Adjustment Method
Let $p_{i2c}=\Pr(X_{i2}=c|x_{i1}^{*}=d,\boldsymbol{z}_{i},w_{i})$ be the working probability that $X_{2}=c$ for unit $i$, given its imputed $x_{i1}^{*}=d$ and $(\boldsymbol{z}_{i},w_{i})$. We determine these from a (multinomial) logistic regression of $X_{2}$ on $X_{1}$ and possibly some function of $\boldsymbol{Z}$ or $W$, with coefficients estimated from the completed data for units with $U_{i}=0$. Note that if we disregard $\boldsymbol{Z}$ and $W$, $p_{i2c}=\Pr(X_{2}=c|X_{1}=d)$ for $c=1,\dots,m_{2}$ and $d=1,\dots,m_{1}$; that is, the conditional probability is the same for all units with $x_{i1}^{*}=d$. For this case (which we use in Section 3.2.2), to simplify notation we drop the index $i$ for individual units and add a subscript $d$ for the value of $X_{1}$, writing $p_{i2c}=p_{2cd}$.
For $c=1,\dots,m_{2}-1$, we sample a plausible value of $\hat{T}_{2c}$ from a $N(T_{2c},V_{2c})$, setting the estimated total for the final level as $\hat{T}_{2m_{2}}=N-\sum_{c=1}^{m_{2}-1}\hat{T}_{2c}$. As in (3.1), we find suitable constants $f_{2c}>0$ to adjust each $p_{i2c}$. We have